OpenStudy (winsteria):

MEDAL AND FAN Which pair of ratios is proportional? A. 4/9 and 16/45 B. 5/12 and 25/144 C. 12/18 and 18/27 D. 15/20 and 7/10

OpenStudy (winsteria):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

so 16/45=16/45 25/144=25/144 18/27=2/3 12/18=2/3 7/10=7/10

OpenStudy (anonymous):

ok see of the division of the 2 pairs gives you a same which is a constant answer then they are proportional

OpenStudy (anonymous):

if* alexander your so fast man 😂

OpenStudy (winsteria):

okay so which one is it? @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

18/27=2/3 12/18=2/3

OpenStudy (winsteria):

Thank you